OpenStudy (anonymous):

How do you add negative numbers? Will medal or fan and please explain @Yungwisdom2000

OpenStudy (anonymous):

@Yungwisdom2000

OpenStudy (anonymous):

ok lets say -2+2=0 ok

OpenStudy (anonymous):

ok

rvc (rvc):

-1+(-1)=?

OpenStudy (anonymous):

I have no idea can you tell me HOW to add them

OpenStudy (anonymous):

that equals 1 its like adding backwards okay heres a rule positive negative= negative positive and positive= positve negative and negative equals negative

rvc (rvc):

+ x - = - - x - = + + x+ = + - x + = - this u have to remember so ans the question @ShaTay

OpenStudy (anonymous):

0... I'm guessing

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

-2 my bad

rvc (rvc):

-1+(-1)=-1-1=-2

OpenStudy (anonymous):

was I right?

rvc (rvc):

good @ShaTay

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

cool thanks guys... @Yungwisdom2000 @rvc

OpenStudy (anonymous):

np

rvc (rvc):

plz... no thanks @ShaTay